OpenStudy (anonymous):

A circle with the radius of 6 in. is placed on a triangle. How do i find the area of the space that the circle doesn't take up?? of, and the radius makes a 90 degree angle with the triangle.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

So you have a triangle with a circle inside it where they intersect at exactly 3 points?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

Just working through a little geometry here, but it seems to me that we have something like this:

OpenStudy (anonymous):

yeah... i think the lines at the top were added in, but thats pretty much what it looks like.

OpenStudy (anonymous):

So that being the case.. If we let x be the length of the hypotenuse of that right triangle, then: \[x = {6 \over sin\ 30^\circ}\]\[Base = x(cos\ 30^\circ) = 6(tan\ 30^\circ)\]\[Height = x + 6 = 6({1 \over sin\ 30^\circ} + 1) \]

OpenStudy (anonymous):

Whoops. The base should be 2 times x(cos 30)

OpenStudy (anonymous):

So 12(tan 30)

OpenStudy (anonymous):

why use trig? we are learning about sectors.... im just wondering a bit.

OpenStudy (anonymous):

Oh, I used trig cause it made the most sense to me. If you know of another way to solve it, feel free.

OpenStudy (anonymous):

oh ok. thanks!!

OpenStudy (anonymous):

I just found the base and height, then can use that to find the area of the triangle. Then I can subtract the area of the circle from that to find the space the circle doesn't occupy.

OpenStudy (anonymous):

i really appreciate your help.... my teacher is didn't show us how to do this kind of problem. and that makes a lot of sense now. thanks!!